OpenStudy (anonymous):

helppppp simplify m=-6 evaluate -m

OpenStudy (anonymous):

Multiply each side by -1.

OpenStudy (anonymous):

so what is -1(-6)

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

no

OpenStudy (anonymous):

dont add

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

but i thought when its a - and a - u add

OpenStudy (anonymous):

-1 times -6 is wht?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yes so-m=

OpenStudy (anonymous):

Think about m = 6. What happens if I subtract 6 from each side? We get m - 6 = 0. Now subtract m from each side. Now -6 = -m. Go through the same process for m = -6.

OpenStudy (anonymous):

so its a 6 or a -6?

OpenStudy (anonymous):

here ill set it up for u

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-m +6=0

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

can u solve it

OpenStudy (anonymous):

m = -6. Add 6 to each side. m + 6 = 0. Now just subtract m from each side to get....

OpenStudy (anonymous):

adam but it says -m

OpenStudy (anonymous):

u guys how can i even subtract m

OpenStudy (anonymous):

like so 6+0 =m

OpenStudy (anonymous):

ohhhhh so 6 =m?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

omg thx

OpenStudy (anonymous):

np

OpenStudy (anonymous):

oh and thats what the other guy said too

OpenStudy (anonymous):

yea ik

OpenStudy (anonymous):

i wish i could give 2 of those lil things

OpenStudy (anonymous):

ill give him 1

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

there we go

OpenStudy (anonymous):

if u need help we will help

OpenStudy (anonymous):

ok :)